HOW many excess electrons must be added to an isolated spherical conductor 0.100 meters in diameter to produce an electric field of 1800N/C just outside the surface?


1
Expert's answer
2021-05-03T10:32:36-0400

The field is


"E=\frac{Q}{4\pi \epsilon_0r^2}=\frac{Ne}{4\pi \epsilon_0r^2},\\\space\\
N=\frac{4\pi \epsilon_0r^2E}{e}=1.25·10^{10}."